Simulate a single-phase AC voltage controller and plot its voltage and current waveforms

using coding in the command window. The plot should be of minimum 300 dpi clarity.

Also include the FFT harmonic spectra of the current.


1
Expert's answer
2021-06-29T00:23:22-0400

Ac voltage controller s a circuit used to convert fixed ac to variable ac without changing the frequency . This conversion with the advent of power electronic devices such as SCR, IGBT, MOSFET, etc. is made very efficiently and flexibly .

The control strategies available to vary fixed ac n ac voltage controller circuits are

  1. On-off cycle control
  2. Firing angle control
  3. Integral cycle switching control


Circuit diagram for simulation is shown below




Input and output wave forms



The expression for rms value of output voltage for resistive load is given by:

"V_o(rms)=V_s \sqrt{\frac{(\pi-\alpha)+(sin2 \alpha)/2}{\pi}}"

where, "V_s=\frac{V_m}{\sqrt{2}}=RMS" Value of input supply voltage


MATLAB circuit simulation diagram




Phase angle control input and output waveforms